# Toward Human-Quantum Computer Interaction: Interface Techniques for Usable Quantum Computing

### v2

@@ 1. Introduction
Quantum computing (QC) offers promising solutions for classically intractable problems that classical computers (those with digital bits) have not been able to solve due to time complexity or memory limitations (Arute2019:quantumSupremacy). For example, factoring a number into two numbers is useful for cryptography, yet classical algorithms typically take exponential time to solve it (days or years) whereas Shor’s algorithm (shors:algorithm) for QC takes only logarithmic time (minutes or days, respectively). Quantum computers are particularly useful in solving problems that requires searching a few valid instances from tons of potential cases, such as drug discovery (cao2018:drug), physics simulation (somma2003:physics), and cryptography (shors:algorithm). With more advances in hardware and software architecture, QC is expected to empower various scientific and engineering areas.
However, using quantum computers requires in-depth knowledge and skills, making it challenging for a wide range of domain experts to adopt QC as a feasible solution. With current tooling like Qiskit (qiskit) or Strawberry Fields (strawberryfields), developers need to write programs by specifying operations for qubits (where a qubit has a value of |$\ket{\psi}=\alpha\ket{0}+\beta\ket{1}$) and read results in terms of bit strings rather than easily-interpretable formats like natural numbers or text. Expert users often need to look at documentation to retrieve detailed information (e.g., physical properties of a machine), yet they are often too basic (e.g., “qubit_properties: Return QubitProperties for a given qubit.”  [^1] https://docs.quantum.ibm.com/api/qiskit/qiskit.providers.BackendV2#qubit_properties) or mathematically oriented, lacking example use cases. QC platforms like IBM and AWS offer graphical tools to explore machines and programs, yet using them requires frequent switching between coding tools and them, and those interfaces often fail to show information with large programs.
Therefore, more intuitive interfaces are highly necessary to mitigate low-level technical hurdles and hence foster QC to users in various domain areas. In response, we, as a team of a design researcher and QC researchers, identify a set of design principles through a survey of existing QC tools and an iterative design process of building QC interfaces. These principles include linking high-level conceptual ideas with low-level quantum information, providing support at different levels of computing, and applying usability standards. Based on these principles, we showcase a set of interaction techniques for usable QC, such as conceptual program writer, hardware dashboard, program viewer, job-sharing API, and problem-specific result viewer. Implementing these techniques as open-sourced high-fidelity prototypes, we demonstrate their usability and technical soundness through case studies that reflect quantum computing learning and research.
@@ 2. Background: Hello Quantum World
We first describe quantum computing (QC) basics, practices, and tools that are relevant to understand this work. We note that the below overview includes simplification for surface-level understanding; refer to Nielsen and Chuang (mike:and:ike) for more precise descriptions. Readers could also use this section as a glossary.
@@ 2.1. Quantum Computing in a Nut Shell
Quantum computers are defined as computing devices that utilize quantum mechanical properties: superposition, entanglement, and interference of quantum particles, as depicted in Figure 1A. The base computation unit of quantum computers is called a qubit, standing for quantum bit, which are manipulated to exhibit those properties. Bits of classical computers (those that we use everyday) are called classical bits or just bits. Superposition describes how quantum states can exist in a probabilistic distribution of discrete measurement outcomes (mike:and:ike). A measurement device measures a qubit by forcing its state into a fixed value from a few discrete values with corresponding probabilities (i.e., non-deterministic). Existing quantum computers typically employ standard measurement that measures a qubit in either |$\ket{0}$ (the ground state) or |$\ket{1}$ (an excited state). Superposition enables the generation of lots of candidate inputs for a given problem, allowing for inherent parallelism during computation. Due to the probabilistic nature of QC, states are often prepared and measured over many trials to form a measurement distribution that represents a quantum application’s outcome.
@@
Next, entanglement describes how two pieces of quantum information must be considered as a composite system after a series of special interactions. Even if two entangled qubits are separated by a great distance, we can infer about the measurement of one by measuring the other (mike:and:ike). For example, suppose two qubits are entangled to have the same state, If a qubit is measured as |$\ket{0}$ (or |$\ket{1}$), then its entangled counterpart is always measured as |$\ket{0}$ (or |$\ket{1}$). Entanglement allows for accessing and manipulating information using its correlation with other information. Lastly, interference means that a manipulation on a qubit affects or cancels previous manipulations on the same qubit. Interference is useful to emphasize desired outcomes or de-emphasize unwanted outcomes.
@@
A quantum computer employs a physical processor to apply superposition, entanglement, and interference to a qubit(s). A high-level classification of quantum computers is whether they manipulate qubits by continuously changing their parameters (analogue) or using discrete gates (digital or gate-based). In this work, we primarily focus on digital quantum computers given its wide availability in practice (e.g., IBM, AWS). For example, a Hadamard gate (or H gate) sets a qubit to be measured as 0 or 1 with a probability 50%, each, implementing the superposition of qubits (Figure 1B). By taking two qubits, a Control-NOT (or C-NOT, CX) gate flips the state of one qubit (|$\ket{0}\rightarrow\ket{1}$, vice versa) only if the other qubit’s state is 1. The state resulting from these operations on two qubits is also called Bell state. An H gate on a qubit and a C-NOT gate on it and another qubit enables the entanglement between them (Figure 1C). By applying multiple consecutive gates on the same qubits, one can realize interference.
@@
A quantum program is typically called a (quantum) circuit, composed of qubits and gates on those qubits. QC developers must be able to deal with this abstraction if they want to debug or create QC algorithms. A quantum circuit is usually represented as in Figure 1C, where each horizontal line represents a qubit and boxes and dots on those lines represent gates applied to the corresponding qubits. The gates in each column in the diagram are those that can be operated at the same time (i.e., in the same layer). After computing all the specified gates, selected qubits are measured as either |$\ket{0}$ or |$\ket{1}$. The measured outcomes are recorded on corresponding classical bits as indicated using the horizontal double-line at the bottom in the circuit diagram. By running the same quantum circuit lots of times, or shots, (typically $1,000$+ shots), one can obtain the probability distribution of outcome states. For example, given a circuit where two qubits $a$ and $b$ are manipulated to a Bell state, the measured outcomes after $1,000$ shots should theoretically have $500$ |$\ket{00}$ states and $500$ |$\ket{11}$ states.
Figure 1. (A) Diagrams for quantum mechanical properties. (B) Example quantum gates. (C) An example quantum circuit that generates a Bell state or quantum entanglement using Hadamard and C-NOT gates on two qubits. Bra ($\bra{}$ |) and ket ($\ket{}$) are a notation for representing quantum states.
@@ 2.2. Quantum Computing Practices
To build and run a quantum program or circuit, a developer typically needs to perform the following tasks: writing a quantum program, selecting a machine(s) to use, optimizing a circuit for the machine to use, running the circuit, and analyzing the output of the program. To help understand, one can make an analogy to data science procedures that commonly include implementing initial statistical or learning models, revising them with different parameters, finding a remote machine if complex computation is needed, and analyzing the computation results like model fits, uncertainty, accuracy, and precision (crisan2021). Below, we overview QC practices based on official tutorials of QC platforms (qiskitGuide; cirqGuide; braketGuide).
@@
Before running the quantum circuit on an actual QC machine (quantum processing unit, QPU), the developer needs to decide which machine to use by considering several factors, such as cost, availability (e.g., shut down, too many pending jobs), architecture, and error constraints. For example, running a QPU on cloud for five minutes can cost a few hundred US dollars, developers need to select simulators first and then use real machines later. Each machine exhibits different gate error rates (i.e., how likely a gate is to cause an error) and decoherence time (i.e., how long a qubit maintains its gate-manipulated state). Given that physical properties must be assessed frequently due to their tenancy to vary over time (dasgupta2021stability), the developer needs to check them before choosing a machine.
After deciding which device to use, the developer needs to optimize or transpile the initial circuit (logical circuit) into a much larger physical circuit. For example, a simple 10-gate logical circuit can result in 500-gate physical circuit. Optimization is necessary because a physical machine only supports a specific set of basis gates which is smaller than that of gates in logical circuits. This is similar to decomposing a wide range of classical logic gates using a smaller set of gates (e.g., an IF classical gate decomposed using NOT and OR classical gates). By merging and decomposing logical gates, optimization also aims to minimize the cumulative gate error rate by reducing the total physical gates and selecting least erroneous qubits and gates. Given exponential possibilities to optimize a logical circuit, many optimization algorithms rely on random seeds and other parameters. We note that this procedure is algorithmic optimization while device-level optimization directly deals with controlling physical devices.
After running a transpiled circuit on a physical machine, the developer must carefully analyze the measurement output. Current, noisy intermediate-scale quantum (NISQ) era machines exhibit non-negligible error rates as hinted earlier, the final output could include some erroneous computations. For example, after running a simple entanglement circuit shown in Figure 1C for 1,000 shots, the theoretical output must include 500 |$\ket{00}$ and 500 |$\ket{11}$ only. However, an actual output may include several |$\ket{01}$ or |$\ket{10}$, and |$\ket{00}$ and |$\ket{11}$ tend to be measured different times.
@@ 2.3. A Survey of Quantum Computing Tools
To provide a snapshot of the state-of-the-art QC tools, we surveyed ten open-sourced libraries for gate-based QC in terms of user interfaces for supporting the above tasks. Specifically, we looked into Qiskit by IBM (qiskit), Cirq by Google (cirq), QuTip (qutip), StrawberryFields (strawberryfields) and PennyLane (pennylane) by Xanadu, Braket by AWS (braket), CUDA-Q by NVIDIA (cudaq), Azure by Microsoft (azureQuantum), qBraid (qbraid), and Classiq (classiq). Our survey is summarized in Table 1.
Circuit composition. Most QC tools offer a Python library for specifying a circuit in a declarative way. Commonly, a developer first need to define a baseline logical circuit in terms of the number of qubits (for operations) and classical bits (for storing measurement outcomes) and then add gates to the circuit. Using Cirq (cirq), for example, a Bell state circuit can be written as circuit.H(0).CNOT(0, 1), where H and CNOT stands for Hadamard and C-NOT gates, and 0 and 1 are the indices for the first two qubits. Azure Quantum (azureQuantum) provides its own programming language, Q#, which needs to define an entry point, similar to C# and JAVA. qBraid (qbraid) and Classiq (classiq) provide graphical interfaces in addition to scripting. qBraid offers a circuit composer (quirk) where a developer can drag and drop gates on a circuit diagram. Since qBraid’s Python library is a wrapper for others like Qiskit and PennyLane, so developers could use expressions from their preferred libraries. Classiq provides an interactive whiteboard (e.g., FigJam, Jamboard), where one can add a post-it for operations and connect them with lines (Figure 2A).
PennyLane and Classiq provide templates for predefined subroutines (e.g., search algorithms, chemical processes) and conceptual representations (e.g., adding numbers, atom representations), reducing the gap between high-level program logic and low-level quantum circuits in part. Yet, developers need to manually change the code for their own use cases. Most tools allow for defining a custom gate by passing a unitary matrix (of which the inverse and conjugate transpose are same). For visual representations, most tools provide internal functions to draw a traditional circuit visualization (e.g., Figure 1C). A few tools support drawing a Bloch sphere (Figure 2B) that show a single qubit’s expected state on a sphere. QuTip supports creating animations for how qubits’ states evolve over operations. Azure supports generative AI-based assistance (Copilot (copilot)) for writing codes for quantum circuits.
Figure 2. Visualizations offered by QC tools. (A) The interactive whiteboard of Classiq (classiq) (B) Bloch sphere. (C) Qiskit’s machine status dashboard (qiskit) (D) Measurement histogram.
@@
Machine selection. In many cases, QC tools let developers filter a machine by its name and provider (the company that owns the machine). Qiskit offers more advanced filters like the number of qubits in a machine, availability, the maximum number of shots, etc. Many QC tools provide separate websites for a dashboard about machines’ status (e.g., Figure 2C). Such dashboards are not directly accessible on an integrated development environment (IDE) like Jupyter Notebook (jupyterNotebooks). While qBraid and Classiq integrate machine status information on their IDE’s, the level of available information is shallow (e.g., not showing error rates). Qiskit and Cirq’s documentations provide templates for how to visualize qubit and operation errors on IDE.
@@
Optimization. Qiskit (qiskit), Cirq (cirq), StrawberryFields (strawberryfields), PennyLane (pennylane), and Classiq (classiq) let developers customize an optimizer by providing parameters like a random seed, the degree of optimization, and use of different compiling algorithms. As optimization procedures typically involve randomization due to exponential possibilities, setting different parameters can result in varying optimization outcomes for the same circuit, which can impact the quality of the program output. Given that optimized circuits share the same data structure with logical circuits, developers usually visualize them for review. Yet, CUDA-Q (cudaq), Braket (braket), and Azure (azureQuantum) incorporated optimization with execution functions, and Classiq (classiq) does not provide optimization results. However, none of these tools provide interfaces for comparing different optimization results; instead, a developer must come up with their own approach.
@@
Results analysis. Most tools offer ways to visualize measurement outcomes (counts) as a histogram (Figure 2D), such as a simple function, code templates as tutorials, and small panels in an IDE. Given that quantum results take a form of bit string (0101…), translating them for problems (e.g., integer, letter, etc.) is a necessary step to interpret measurement outcomes StrawberryFields and PennyLane’s tutorials offer several ways to do so, such as decision boundaries [^4] https://pennylane.ai/qml/demos/tutorial_kernels_module/. While measurement outcomes involve some errors, related uncertainty information is not usually communicated.
@@
Takeaways. We summarize a few takeaways from this survey with respect to common usability support. (1) QC libraries provide raw visualization and circuit codes as templates in their tutorials rather than application programming interfaces (APIs) integrated to themselves. Because QC users are not necessarily classical programming experts and it is challenging to edit raw codes (mei2018:vistoolsurvey; battle2022:d3), more integrated interface support for frequently used visualizations seems to be useful. (2) Useful information is often not accessible on IDE. QC developers often need to switch platform’s dashboard, IDE, documentations, and tutorials, which can make their tasks more tedious than needed. (3) It is hard to make interactions for on-demand details. Visualizations templates and APIs are often designed in a static way, limited to support chained interactions for seeking details. For example, static circuit visualization for large circuits can be confusing to identify the relationship between qubits and gates. To fill this gap, our work aims to find methods to integrate this kind of usability support for QC tools.

### v3

@@ 1. Introduction
Quantum computing (QC) offers promising solutions for classically intractable problems that classical computers (those with digital bits) have not been able to solve due to time complexity or memory limitations (Arute et al., 2019). For example, factoring a number into two numbers is useful for cryptography, yet classical algorithms typically take exponential time to solve it (days or years) whereas Shor’s algorithm (Shor, 1994) for QC takes only logarithmic time (minutes or days, respectively). Quantum computers are particularly useful in solving problems that requires searching a few valid instances from tons of potential cases, such as drug discovery (Cao et al., 2018), physics simulation (Somma et al., 2003), and cryptography (Shor, 1994). With more advances in hardware and software architecture, QC is expected to empower various scientific and engineering areas.
However, using quantum computers requires in-depth knowledge and skills, making it challenging for a wide range of domain experts to adopt QC as a feasible solution. With current tooling like Qiskit (IBM, 2017) or Strawberry Fields (Xanadu, 2018), developers need to write programs by specifying operations for qubits (where a qubit has a value of $|{\psi}\rangle=\alpha|{0}\rangle+\beta|{1}\rangle$) and read results in terms of bit strings rather than easily-interpretable formats like natural numbers or text. Expert users often need to look at documentation to retrieve detailed information (e.g., physical properties of a machine), yet they are often too basic (e.g., “qubit_properties: Return QubitProperties for a given qubit.”  [^1] https://docs.quantum.ibm.com/api/qiskit/qiskit.providers.BackendV2#qubit_properties) or mathematically oriented, lacking example use cases. QC platforms like IBM and AWS offer graphical tools to explore machines and programs, yet using them requires frequent switching between coding tools and them, and those interfaces often fail to show information with large programs.
Therefore, more intuitive interfaces are highly necessary to mitigate low-level technical hurdles and hence foster QC to users in various domain areas. In response, we, as a team of a design researcher and QC researchers, identify a set of design principles through a survey of existing QC tools and an iterative design process of building QC interfaces. These principles include linking high-level conceptual ideas with low-level quantum information, providing support at different levels of computing, and applying usability standards. Based on these principles, we showcase a set of interaction techniques for usable QC, such as conceptual program writer, hardware dashboard, program viewer, job-sharing API, and problem-specific result viewer. Implementing these techniques as open-sourced high-fidelity prototypes, we demonstrate their usability and technical soundness through use cases that reflect QC learning and research.
@@ 2. Background: Hello Quantum World
We first describe quantum computing (QC) basics, practices, and tools that are relevant to understand this work. We note that the below overview includes simplification for surface-level understanding; refer to Nielsen and Chuang (Nielsen and Chuang, 2010) for more precise descriptions. Readers could also use this section as a glossary.
@@ 2.1. Quantum Computing in a Nut Shell
Quantum computers are defined as computing devices that utilize quantum mechanical properties: superposition, entanglement, and interference of quantum particles, as depicted in Figure 1A. The base computation unit of quantum computers is called a qubit, standing for quantum bit, which are manipulated to exhibit those properties. Bits of classical computers (those that we use everyday) are called classical bits or just bits. Superposition describes how quantum states can exist in a probabilistic distribution of discrete measurement outcomes (Nielsen and Chuang, 2010). A measurement device measures a qubit by forcing its state into a fixed value from a few discrete values with corresponding probabilities (i.e., non-deterministic). Existing quantum computers typically employ standard measurement that measures a qubit in either $|{0}\rangle$ (the ground state) or $|{1}\rangle$ (an excited state). Superposition enables the generation of lots of candidate inputs for a given problem, allowing for inherent parallelism during computation. Due to the probabilistic nature of QC, states are often prepared and measured over many trials to form a measurement distribution that represents a quantum application’s outcome.
@@
Next, entanglement describes how two pieces of quantum information must be considered as a composite system after a series of special interactions. Even if two entangled qubits are separated by a great distance, we can infer about the measurement of one by measuring the other (Nielsen and Chuang, 2010). For example, suppose two qubits are entangled to have the same state. If a qubit is measured as $|{0}\rangle$ (or $|{1}\rangle$), then its entangled counterpart is always measured as $|{0}\rangle$ (or $|{1}\rangle$). Entanglement allows for accessing and manipulating information using its correlation with other information. Lastly, interference means that a manipulation on a qubit affects or cancels previous manipulations on the same qubit. Interference is useful to emphasize desired outcomes or de-emphasize unwanted outcomes.
@@
A quantum computer employs a physical processor to apply superposition, entanglement, and interference to a qubit(s). A high-level classification of quantum computers is whether they manipulate qubits by continuously changing their parameters (analogue) or using discrete gates (digital or gate-based). In this work, we primarily focus on digital quantum computers given its wide availability in practice (e.g., IBM, AWS). For example, a Hadamard gate (or H gate) sets a qubit to be measured as 0 or 1 with a probability of 50%, each, implementing the superposition of qubits (Figure 1B). By taking two qubits, a Control-NOT (or C-NOT, CX) gate flips the state of one qubit ($|{0}\rangle\rightarrow|{1}\rangle$, vice versa) only if the other qubit’s state is $|{1}\rangle$. The state resulting from these operations on two qubits is also called Bell state. An H gate on a qubit and a C-NOT gate on it and another qubit enables the entanglement between them (Figure 1C). By applying multiple consecutive gates on the same qubits, one can realize interference.
@@
A quantum program is typically called a (quantum) circuit, composed of qubits and gates on those qubits. QC developers must be able to deal with this abstraction if they want to debug or create QC algorithms. A quantum circuit is usually represented as in Figure 1C, where each horizontal line represents a qubit and boxes and dots on those lines represent gates applied to the corresponding qubits. The gates in each column in the diagram are those that can be operated at the same time (i.e., in the same layer). After computing all the specified gates, selected qubits are measured as either $|{0}\rangle$ or $|{1}\rangle$. The measured outcomes are recorded on corresponding classical bits as indicated using the horizontal double-line at the bottom in the circuit diagram. By running the same quantum circuit lots of times, or shots, (typically $1,000$+ shots), one can obtain the probability distribution of outcome states. For example, given a circuit where two qubits $a$ and $b$ are manipulated to a Bell state, the measured outcomes after $1,000$ shots should theoretically have $500$ $|{00}\rangle$ states and $500$ $|{11}\rangle$ states.
[…]
@@ 2.2. Quantum Computing Practices
To build and run a quantum program or circuit, a developer typically needs to perform the following tasks: writing a quantum program, selecting a machine(s) to use, optimizing a circuit for the machine to use, running the circuit, and analyzing the output of the program. To help understand, one can make an analogy to data science procedures that commonly include implementing initial statistical or learning models, revising them with different parameters, finding a remote machine if complex computation is needed, and analyzing the computation results like model fits, uncertainty, accuracy, and precision (Crisan et al., 2021). Below, we overview QC practices based on official tutorials of QC platforms (IBM, 2024; Google, 2024; AWS, 2024).
@@
Before running the quantum circuit on an actual QC machine (quantum processing unit, QPU), the developer needs to decide which machine to use by considering several factors, such as cost, availability (e.g., shut down, too many pending jobs), architecture, and error constraints. For example, running a QPU on cloud for five minutes can cost a few hundred US dollars, developers need to select simulators first and then use real machines later. Each machine exhibits different gate error rates (i.e., how likely a gate is to cause an error) and decoherence time (i.e., how long a qubit maintains its gate-manipulated state). Given that physical properties must be assessed frequently due to their tenancy to vary over time (Dasgupta and Humble, 2021), the developer needs to check them before choosing a machine.
After deciding which device to use, the developer needs to optimize or transpile the initial circuit (logical circuit) into a much larger physical circuit. For example, a simple 10-gate logical circuit can result in a 500-gate physical circuit. Optimization is necessary because a physical machine only supports a specific set of basis gates which is smaller than that of gates in logical circuits. This is similar to decomposing a wide range of classical logic gates using a smaller set of gates (e.g., an IF classical gate decomposed using NOT and OR classical gates). By merging and decomposing logical gates, optimization also aims to minimize the cumulative gate error rate by reducing the total physical gates and selecting least erroneous qubits and gates. Given exponential possibilities to optimize a logical circuit, many optimization algorithms rely on random seeds and other parameters. We note that this procedure is algorithmic optimization while device-level optimization directly deals with controlling physical devices.
After running a transpiled circuit on a physical machine, the developer must carefully analyze the measurement output. Current, noisy intermediate-scale quantum (NISQ) era machines exhibit non-negligible error rates as hinted earlier, the final output could include some erroneous computations. For example, after running a simple entanglement circuit shown in Figure 1C for 1,000 shots, the theoretical output must include 500 $|{00}\rangle$ and 500 $|{11}\rangle$ only. However, an actual output may include several $|{01}\rangle$ or $|{10}\rangle$, and $|{00}\rangle$ and $|{11}\rangle$ tend to be measured different times.
@@ 2.3. A Survey of Quantum Computing Tools
To provide a snapshot of the state-of-the-art QC tools, we surveyed ten open-sourced libraries for gate-based QC in terms of user interfaces for supporting the above tasks. Specifically, we looked into Qiskit by IBM (IBM, 2017), Cirq by Google (Google, 2018), QuTip (QuTip, 2011), StrawberryFields (Xanadu, 2018) and PennyLane (Xanadu, 2021) by Xanadu, Braket by AWS (AWS, 2019), CUDA-Q by NVIDIA (NVDIA, 2023), Azure by Microsoft (Microsoft, 2023), qBraid (qBraid, 2022), and Classiq (Classiq, 2024). Our survey is summarized in Table 1.
Circuit composition. Most QC tools offer a Python library for specifying a circuit in a declarative way. Commonly, a developer first need to define a baseline logical circuit in terms of the number of qubits (for operations) and classical bits (for storing measurement outcomes) and then add gates to the circuit. Using Cirq (Google, 2018), for example, a Bell state circuit can be written as circuit.H(0).CNOT(0, 1), where H and CNOT stands for Hadamard and C-NOT gates, and 0 and 1 are the indices for the first two qubits. Azure Quantum (Microsoft, 2023) provides its own programming language, Q#, which needs to define an entry point, similar to C# and JAVA. qBraid (qBraid, 2022) and Classiq (Classiq, 2024) provide graphical interfaces in addition to scripting. qBraid offers a circuit composer (Gidney, 2016) where a developer can drag and drop gates on a circuit diagram. Since qBraid’s Python library is a wrapper for others like Qiskit and PennyLane, so developers could use expressions from their preferred libraries. Classiq provides an interactive whiteboard (e.g., FigJam, Jamboard), where one can add a post-it for operations and connect them with lines (Figure 2A).
PennyLane and Classiq provide templates for predefined subroutines (e.g., search algorithms, chemical processes) and conceptual representations (e.g., adding numbers, atom representations), reducing the gap between high-level program logic and low-level quantum circuits in part. Yet, developers need to manually change the code for their own use cases. Most tools allow for defining a custom gate by passing a unitary matrix (of which the inverse and conjugate transposes are same). For visual representations, most tools provide internal functions to draw a traditional circuit visualization (e.g., Figure 1C). A few tools support drawing a Bloch sphere (Figure 2B) that show a single qubit’s expected state on a sphere. QuTip supports creating animations for how qubits’ states evolve over operations. Azure supports generative AI-based assistance (Copilot (Microsoft, 2024)) for writing codes for quantum circuits.
Figure 2. Visualizations offered by QC tools. (A) The interactive whiteboard of Classiq (Classiq, 2024) (B) Bloch sphere. (C) Qiskit’s machine status dashboard (IBM, 2017) (D) Measurement histogram.
@@
Machine selection. In many cases, QC tools let developers filter a machine by its name and provider (the company that owns the machine). Qiskit offers more advanced filters like the number of qubits in a machine, availability, the maximum number of shots, etc. Many QC tools provide separate websites for a dashboard about machines’ status (e.g., Figure 2C). Such dashboards are not directly accessible on an integrated development environment (IDE) like Jupyter Notebook (Kluyver et al., 2016). While qBraid and Classiq integrate machine status information on their IDE’s, the level of available information is shallow (e.g., not showing error rates). Qiskit and Cirq’s documentations provide templates for how to visualize qubit and operation errors on IDE.
@@
Optimization. Qiskit (IBM, 2017), Cirq (Google, 2018), StrawberryFields (Xanadu, 2018), PennyLane (Xanadu, 2021), and Classiq (Classiq, 2024) let developers customize an optimizer by providing parameters like a random seed, the degree of optimization, and the use of different algorithms. As optimization procedures typically involve randomization due to exponential possibilities, setting different parameters can result in varying optimization outcomes for the same circuit, which can impact the quality of the program output. Given that optimized circuits share the same data structure with logical circuits, developers usually visualize them for review. Yet, CUDA-Q (NVDIA, 2023), Braket (AWS, 2019), and Azure (Microsoft, 2023) incorporated optimization with execution functions, and Classiq (Classiq, 2024) does not provide optimization results. However, none of these tools provide interfaces for comparing different optimization results; instead, a developer must come up with their own approach.
@@
Results analysis. Most tools offer ways to visualize measurement outcomes (counts) as a histogram (Figure 2D), such as a simple function, code templates as tutorials, and small panels in an IDE. Given that quantum results take a form of bit string (0101…), translating them for problems (e.g., integer, letter, etc.) is a necessary step to interpret measurement outcomes. StrawberryFields and PennyLane’s tutorials offer several ways to do so, such as decision boundaries [^4] https://pennylane.ai/qml/demos/tutorial_kernels_module/. While measurement outcomes involve some errors, related uncertainty information is not usually communicated.
@@
Takeaways. We summarize a few takeaways from this survey with respect to common usability support. (1) QC libraries provide raw visualization and circuit codes as templates in their tutorials rather than application programming interfaces (APIs) integrated to themselves. Because QC users are not necessarily classical programming experts and it is challenging to edit raw codes (Mei et al., 2018; Battle et al., 2022), more integrated interface support for frequently used visualizations seems to be useful. (2) Useful information is often not accessible on IDE. QC developers often need to switch platform’s dashboard, IDE, documentations, and tutorials, which can make their tasks more tedious than needed. (3) It is hard to make interactions for on-demand details. Visualization templates and APIs are often designed in a static way, limited to support chained interactions for seeking details. For example, static circuit visualization for large circuits can be confusing to identify the relationship between qubits and gates. To fill this gap, our work aims to find methods to integrate this kind of usability support for QC tools.